OpenStudy (anonymous):

f(x) =square root of (1-x) power of 3 over square root three of x it is the function even , odd or neither even nor odd , explain :)

OpenStudy (amistre64):

test f(a) and f(-a) if f(a) = f(-a) ; its even if f(a) = -f(a) ; its odd if neither, then its neither

OpenStudy (anonymous):

i try but still i can't get the answer

OpenStudy (anonymous):

can you try to type the solution , thank you

OpenStudy (amistre64):

f(a) = sqrt(1-a)^3 ---------- 3a f(-a) = sqrt(1+a)^3 ---------- -3a i got no idea if im reading your grammartized stuff correctly tho

OpenStudy (amistre64):

1-1 = 0 these aint gonna match up, but its never a good idea to go off of 1 1+1 = 2 example my gut says neither, but without knowng what the problem really looks like, I cant be sure

OpenStudy (amistre64):

yeah, my interpretation of your stuff has no symmetry about anything :) http://www.wolframalpha.com/input/?i=sqrt%281-x%29%5E3%2F3x